OpenStudy (anonymous):

\[\int\limits_{1}^{In3}(e^x-e^{3x})/(1+e^x)\]

OpenStudy (anonymous):

\[\int\limits_{1}^{In3}(e^x-e^{3x})/(1+e^x)\]

OpenStudy (anonymous):

simplify..

OpenStudy (anonymous):

e^x (1+e^x)=e^x+e^2x

OpenStudy (anonymous):

its o not e^x+e^2x, but e^x-e^3

OpenStudy (anonymous):

if e^x = t and e^xdx = dt you got: \[(1-t ^{2})dt/1+t = 1-t dt\] rest it's easy

OpenStudy (anonymous):

 now the integral should be easier...